# Vision-Centric Activation and Coordination for Multimodal Large Language Models

### v1

@@ 1 Introduction
Benefiting from advancements in LLMs OpenAI (2022); Touvron et al. (2023), MLLMs have exhibited remarkable capabilities in various visual understanding tasks, such as visual question answering Achiam et al. (2023); Liu et al. (2023), object grounding Wang et al. (2023a), and referring expression comprehension You et al. (2023). These innovative models typically incorporate images as multimodal tokens into LLMs, utilizing features from pre-trained vision-language models (e.g., CLIP Radford et al. (2021)) to facilitate language-output understanding. Within this architecture, although the effectiveness of advanced LLMs in enhancing multimodal capabilities is well-established, the potential of visual components remains insufficiently underscored, posing a bottleneck to achieving comprehensive visual understanding Wang et al. (2024a); Tong et al. (2024a).
@@
Figure 1: Visualization of Visual Activation Methods: (a) Attention Visualization. Utilizing samples from MMVP Tong et al. (2024b), we investigate which visual regions dominate the final predictions of MLLMs. Specifically, since MLLMs are anticipated to deliver answers at the final position of the input sequence, we visualize the attention scores between the last token and all image tokens. Compared to the other schemes (i.e., (i) LLaVA Liu et al. (2023), (iii) multi-encoder scheme Tong et al. (2024a), (iv) reconstructive scheme Wang et al. (2024a) and (iv) the proposed VaCo), our vision-centric activation effectively directs MLLMs to focus on the crucial visual areas mentioned in the questions (Q) (e.g., “palm” and “fingers” in the first example), leading to correct answers; (b) Inference Example. VaCo facilitates MLLM in delivering comprehensive answers (A) through the assistance of diverse Visual Results. Note that visual outputs are an optional byproduct.
@@
To address these issues, we propose VaCo, a tuning approach for MLLMs that leverages Vision-Centric activation and Coordination via existing VFMs, facilitating the vision-understanding capabilities of MLLMs, as illustrated in Figure 1 (a). Specifically, we advocate providing supplementary query-based visual supervision, emphasizing discriminative alignment objectives across various VFMs to preserve perceptual cues, as opposed to the generative objective adopted in recent work Wang et al. (2024a). The primary challenge revolves around converting MLLM representations into distinct task-aware spaces, facilitating the activation of comprehensive visual knowledge under the supervision of multiple VFMs. To this end, we design learnable Modular Task Queries (MTQs) (e.g., depth queries) for specific VFM and feed them together with the vision token into the MLLM. These MTQs aim to extract task-specific visual information from the visual representations within MLLMs. Furthermore, we introduce Visual Alignment Layers (VALs), which convert MTQs into Visual Alignment Queries (VAQs) that are dimensionally aligned with a specific vision task. Through the collaboration of modular task and visual alignment queries, the causally structured MLLM is instructed to first output MTQs encapsulating specific perceptual priors before generating the final answer, as presented in Figure 1 (b). Thus, the query-based VaCo indirectly activates fine-grained vision information within the original MLLM representations. Although we employ separate queries for different VFMs, representational conflicts Metz et al. (2020); Vandenhende et al. (2021) still potentially arise among multiple groups of MTQs processed by the causal MLLM. Therefore, we present the Token Gateway Mask (TGM) to restrict the information flow based on MTQ type, coordinating the diverse representations across multiple VFMs. As shown in Figure 2, VaCo promotes the convergence of the next-token prediction for textual tokens, showcasing its effectiveness in activating and coordinating visual priors in the two-stage MLLM training framework Liu et al. (2023).
@@ 2 Preliminary
Multimodal Large Language Models. A LLM Radford et al. (2018; 2019) parameterized by $\theta$ aims to model the canonical causal distribution $p_{\theta}(\bm{x})$ over a sequence of text tokens $\{\bm{x}_{i}\}_{i=1}^{T}$ with length $T$, denoted as $p_{\theta}(\bm{x})=\prod_{t=1}^{T}p_{\theta}(\bm{x}\mid\bm{x}_{<t})$. To achieve effective multimodal understanding, MLLMs Liu et al. (2023) expand the original text tokens with $K$-length visual tokens $\{\bm{v}_{i}\}_{k=1}^{K}$ as the input sequences of LLM. A typical method to obtain these visual tokens involves processing image $\bm{I}$ through the patch embedding of a $\zeta$-parameterized visual encoder $\mathcal{E}_{\zeta}$, such as CLIP Radford et al. (2021) and SigLIP Zhai et al. (2023). Subsequently, a $\phi$-parameterized multimodal projector $\mathcal{P}_{\phi}$ is employed to map these visual tokens into the feature space of LLMs. Then, the maximum likelihood estimation (MLE) objective function of an MLLM can be formulated as follows:
@@ 3.1 Vision-Centric Activation
where $\mathcal{F}$ denotes the pre-trained VFM, which remains frozen during the tuning process, $B$ denotes the number of samples, $\lambda=0.1$ is the balancing factor, and $\tau$ denotes the learnable scaling factor for the contrastive item. Through the optimization of this objective, coupled with the synergy between MTQs and VAQs, we activate the existing visual representation of MLLM by indirectly incorporating the expert knowledge from purely visual models without text-visual alignment. Compared to existing multi-encoder schemes Tong et al. (2024a); Jain et al. (2024), our method relies solely on MTQs during inference without additional visual encoders. Therefore, VaCo efficiently utilizes diverse visual cues while significantly reducing computational costs.
@@ 3.3 Optimization Objective
Given the MTQs, textual responses from MLLMs, and vision-centric targets from VFMs, we tune the image projector, MTQs, VAL, and LLM following a two-stage training recipe Liu et al. (2023), i.e., the pre-training (PT) stage and the supervised fine-tuning (SFT) stage. The training loss of the MLLM equipped with the proposed MTQs, VALs, and TGM can be summarized as follows:
@@ 4 Experiments
Implementation Details. We conduct all experiments based on the LLaVA architecture. In particular, we utilize Vicuna-1.5-7B Chiang et al. (2023) and Qwen2-7B Bai et al. (2023) as backbones, alongside the CLIP-ViT-L/14@336 Radford et al. (2021) and SigLIP-ViT-SO400M/14@384 Zhai et al. (2023) as visual encoders, to assess the effectiveness of our proposed approach. In the VaCo framework, we employ DINOv2 Oquab et al. (2023), OneFormer Jain et al. (2023), Depth Anything V2 Yang et al. (2024a), and VGGT Wang et al. (2025) as the VFMs. For the pre-training and supervised fine-tuning phases, we employ the LLavA-558K Liu et al. (2023) and Cambrian-737K Tong et al. (2024a) datasets, respectively. To demonstrate the improvement of understanding ability at the region/scene-level, we integrate the AS-V2 datasets Wang et al. (2024b), which annotates the question-answering pairs in the formats of grounding and relation. We perform the pre-training phase for $1$ epoch, utilizing a learning rate of $2e^{-3}$ and a batch size of $256$. We fine-tune the model for $3$ epochs, with a learning rate of $2e^{-5}$ and a batch size of $128$. The length $Q$ of MTQs for a single task is set to $8$. We optimize both phases with the AdamW optimizer Kingma & Ba (2014). We implement all experiments on 8 NVIDIA H20 GPUs, each with $98$ GB of memory. We carry out the majority of metric evaluations using the VLEMEval library Duan et al. (2024), except when certain metrics are not covered.
@@ 4.1 Quantitative Comparison
Results on General Benchmarks. To assess the general capabilities of our VaCo, we conduct a comprehensive comparison with prominent MLLMs, as presented in Table 3.3. The benchmarks encompass diverse sub-tasks that evaluate various fine-grained capabilities, such as logic/attribute/relation understanding and coarse/fine-grained perception in MMBench Liu et al. (2024b). Benefiting from the stronger visual comprehension ability provided by vision-centric activation and coordination, VaCo achieves superior performance on these benchmarks. For instance, our VaCo attains an overall score of $78.6$ on MMBench-English and $46.7$ on MMMU Yue et al. (2024), outperforming LLaVA-1.5 Liu et al. (2023) by $2.3$ points and $4.9$ points, respectively. Notably, compared to ROSS Wang et al. (2024a) that employs intrinsic visual activation with a reconstructive objective, our VaCo demonstrates significant advantages on benchmarks such as SEED Li et al. (2023a), MMVP Tong et al. (2024b), RealWorld, and AI2D Hiippala et al. (2021). This indicates that incorporating a query-based discriminative distillation objective with perceptual priors across multiple visual foundation models is more effective. Besides, our VaCo allows MLLMs to utilize a single visual encoder during inference, avoiding additional computational overhead from multiple encoders. Despite this, VaCo achieves superior performance across various metrics compared to Cambrian-1 Tong et al. (2024a), which aggregates multiple visual encoders (CLIP Radford et al. (2021), SigLIP Zhai et al. (2023), DINOv2 Oquab et al. (2023), and ConvNext Liu et al. (2022)). These results highlight the efficacy of vision-centric activation and coordination in improving the general visual-language understanding capabilities of MLLMs.
@@ 4.2 Ablation Study
In Table 4, we investigate the effects of various visual activation strategies and VFMs. We adopt LLaVA-v1.5 Liu et al. (2023) baseline for all ablation studies, which integrates Vicuna-7B as the backbone and CLIP-ViT-L/14@336 as the vision encoder. We train our VaCo on LLavA-558K and Cambrian-737K datasets during the pre-training and the instruction tuning stages, respectively. All variants of VaCo are evaluated on (1) HaluBench Guan et al. (2024): Hall.; (2) MMBEN: MMBench-English Liu et al. (2024b); (3) SEEDI: SEED-Image Li et al. (2023a); (4) MMMU Yue et al. (2024); (5) MMVP Tong et al. (2024b); and (6) Real.: RealWorldQA x.ai (2024) benchmarks.
@@ 5.1 Multimodal Large Language Models
Recent advances in LLMs Achiam et al. (2023); Chiang et al. (2023); Touvron et al. (2023); Zeng et al. (2022) have driven substantial progress in multimodal LLMs (MLLMs) for visual recognition and understanding. LLM-based multimodal models Liu et al. (2023); Bai et al. (2023); Chen et al. (2023a; 2024); Wang et al. (2023b); Zhai et al. (2022); Zhu et al. (2023) aim to integrate the advanced understanding capabilities of LLMs with the visual information provided by language-supervised visual encoders Radford et al. (2021); Zellers et al. (2019); Zhai et al. (2023); Tong et al. (2024b). A connector Liu et al. (2023; 2024a); Alayrac et al. (2022); Li et al. (2023b); Ge et al. (2023); Li et al. (2024a) is employed to convert visual signals from the visual encoder into visual tokens within the LLM representational space. Despite these models exhibiting powerful performance, the lack of explicit visual supervision inherently leads to the degradation of rich visual information. Recent studies Wang et al. (2024a); Huang et al. (2025) have emphasized that activating visual signals boosts the comprehension capabilities of MLLMs. In contrast to ROSS Wang et al. (2024a), which reconstructs single visual features from MLLM visual tokens, our VaCo adopts a query-based pipeline that enables more flexible alignment of task-aware perceptual priors across multiple VFMs.
@@ 5.2 Vision Foundation Models in MLLMs
The pioneering LLaVA Liu et al. (2023) model adopts CLIP Radford et al. (2021) as its visual encoder, and subsequent work has integrated stronger vision–language alignment encoders Zhai et al. (2023); Chen et al. (2024) into MLLM architectures. Building upon these approaches, some models Tong et al. (2024a; b) incorporate multiple visual encoders, such as DINOv2 Oquab et al. (2023), aMAE He et al. (2022) and MoCoV3 He et al. (2020), alongside vision-language alignment models. In recent years, certain works Jain et al. (2024); Huang et al. (2025) have discovered that visual signals from task-specific models address the limitation of CLIP, which mainly captures global information. Consequently, task-specific models such as DepthAnything Yang et al. (2024a; b), OneFormer Jain et al. (2023), FLARE Zhang et al. (2025) and VGGT Wang et al. (2025) have potential for comprehensive integration into MLLMs. Nonetheless, multi-encoder schemes, such as Cambrain-1 Tong et al. (2024a) and VCoder Jain et al. (2024), directly feed multi-source features into MLLMs, increasing the visual encoding cost while leaving potential representation conflicts unaddressed. Unlike these works, our VaCo leverages learnable task-specific queries to activate and coordinate the knowledge from multiple VFMs, thereby enhancing the visual comprehension capabilities of MLLMs without significantly increasing computational overhead.

### v3

@@ 1 Introduction
Benefiting from advancements in LLMs OpenAI (2022); Touvron et al. (2023), MLLMs have exhibited remarkable capabilities in various visual understanding tasks, such as visual question answering Achiam et al. (2023); Liu et al. (2023a), object grounding Wang et al. (2023a), and referring expression comprehension You et al. (2023). These innovative models typically incorporate images as multimodal tokens into LLMs, utilizing features from pre-trained vision-language models (e.g., CLIP Radford et al. (2021)) to facilitate language-output understanding. Within this architecture, although the effectiveness of advanced LLMs in enhancing multimodal capabilities is well-established, the potential of visual components remains insufficiently underscored, posing a bottleneck to achieving comprehensive visual understanding Wang et al. (2024a); Tong et al. (2024a).
@@
Figure 1: Visualization of Visual Activation Methods: (a) Attention Visualization. Utilizing samples from MMVP Tong et al. (2024b), we investigate which visual regions dominate the final predictions of MLLMs. Specifically, since MLLMs are anticipated to deliver answers at the final position of the input sequence, we visualize the attention scores between the last token and all image tokens. Compared to the other schemes (i.e., (i) LLaVA Liu et al. (2023a), (iii) multi-encoder scheme Tong et al. (2024a), (iv) reconstructive scheme Wang et al. (2024a) and (iv) the proposed VaCo), our vision-centric activation effectively directs MLLMs to focus on the crucial visual areas mentioned in the questions (Q) (e.g., “palm” and “fingers” in the first example), leading to correct answers; (b) Inference Example. VaCo facilitates MLLM in delivering comprehensive answers (A) through the assistance of diverse Visual Results. Note that visual outputs are an optional byproduct.
@@
To address these issues, we propose VaCo, a tuning approach for MLLMs that leverages Vision-Centric activation and Coordination via existing VFMs, facilitating the vision-understanding capabilities of MLLMs, as illustrated in Figure 1 (a). Specifically, we advocate providing supplementary query-based visual supervision, emphasizing discriminative alignment objectives across various VFMs to preserve perceptual cues, as opposed to the generative objective adopted in recent work Wang et al. (2024a). The primary challenge revolves around converting MLLM representations into distinct task-aware spaces, facilitating the activation of comprehensive visual knowledge under the supervision of multiple VFMs. To this end, we design learnable Modular Task Queries (MTQs) (e.g., depth queries) for specific VFM and feed them together with the vision token into the MLLM. These MTQs aim to extract task-specific visual information from the visual representations within MLLMs. Furthermore, we introduce Visual Alignment Layers (VALs), which convert MTQs into Visual Alignment Queries (VAQs) that are dimensionally aligned with a specific vision task. Through the collaboration of modular task and visual alignment queries, the causally structured MLLM is instructed to first output MTQs encapsulating specific perceptual priors before generating the final answer, as presented in Figure 1 (b). Thus, the query-based VaCo indirectly activates fine-grained vision information within the original MLLM representations. Although we employ separate queries for different VFMs, representational conflicts Metz et al. (2020); Vandenhende et al. (2021) still potentially arise among multiple groups of MTQs processed by the causal MLLM. Therefore, we present the Token Gateway Mask (TGM) to restrict the information flow based on MTQ type, coordinating the diverse representations across multiple VFMs. As shown in Figure 2, VaCo promotes the convergence of the next-token prediction for textual tokens, showcasing its effectiveness in activating and coordinating visual priors in the two-stage MLLM training framework Liu et al. (2023a).
@@ 2 Preliminary
Multimodal Large Language Models. A LLM Radford et al. (2018; 2019) parameterized by $\theta$ aims to model the canonical causal distribution $p_{\theta}(\bm{x})$ over a sequence of text tokens $\{\bm{x}_{i}\}_{i=1}^{T}$ with length $T$, denoted as $p_{\theta}(\bm{x})=\prod_{t=1}^{T}p_{\theta}(\bm{x}\mid\bm{x}_{<t})$. To achieve effective multimodal understanding, MLLMs Liu et al. (2023a) expand the original text tokens with $K$-length visual tokens $\{\bm{v}_{i}\}_{k=1}^{K}$ as the input sequences of LLM. A typical method to obtain these visual tokens involves processing image $\bm{I}$ through the patch embedding of a $\zeta$-parameterized visual encoder $\mathcal{E}_{\zeta}$, such as CLIP Radford et al. (2021) and SigLIP Zhai et al. (2023). Subsequently, a $\phi$-parameterized multimodal projector $\mathcal{P}_{\phi}$ is employed to map these visual tokens into the feature space of LLMs. Then, the maximum likelihood estimation (MLE) objective function of an MLLM can be formulated as follows:
@@ 3.1 Vision-Centric Activation
where $\mathcal{F}$ denotes the pre-trained VFM, which remains frozen during the tuning process, $B$ denotes the number of samples, $\lambda=0.1$ is the balancing factor, and $\tau$ denotes the learnable scaling factor for the contrastive item. Through the optimization of this objective, coupled with the synergy between MTQs and VAQs, we activate the existing visual representation of MLLM by indirectly incorporating the expert knowledge from purely visual models without text-visual alignment. Compared to existing multi-encoder schemes Tong et al. (2024a); Jain et al. (2024), our method relies solely on MTQs during inference without additional visual encoders. Therefore, VaCo efficiently utilizes diverse visual cues while significantly reducing computational costs (see Table 8 of Appendix).
@@ 3.3 Optimization Objective
Given the MTQs, textual responses from MLLMs, and vision-centric targets from VFMs, we tune the image projector, MTQs, VAL, and LLM following a two-stage training recipe Liu et al. (2023a), i.e., the pre-training (PT) stage and the supervised fine-tuning (SFT) stage. The training loss of the MLLM equipped with the proposed MTQs, VALs, and TGM can be summarized as follows:
@@ 4 Experiments
Implementation Details. We conduct all experiments based on the LLaVA architecture. In particular, we utilize Vicuna-1.5-7B Chiang et al. (2023) and Qwen2-7B Bai et al. (2023) as backbones, alongside the CLIP-ViT-L/14@336 Radford et al. (2021) and SigLIP-ViT-SO400M/14@384 Zhai et al. (2023) as visual encoders, to assess the effectiveness of our proposed approach. In the VaCo framework, we employ DINOv2 Oquab et al. (2023), OneFormer Jain et al. (2023), Depth Anything V2 Yang et al. (2024a), and VGGT Wang et al. (2025) as the VFMs. For the pre-training and supervised fine-tuning phases, we employ the LLavA-558K Liu et al. (2023a) and Cambrian-737K Tong et al. (2024a) datasets, respectively. To demonstrate the improvement of understanding ability at the region/scene-level, we integrate the AS-V2 datasets Wang et al. (2024b), which annotates the question-answering pairs in the formats of grounding and relation. We perform the pre-training phase for $1$ epoch, utilizing a learning rate of $2e^{-3}$ and a batch size of $256$. We fine-tune the model for $3$ epochs, with a learning rate of $2e^{-5}$ and a batch size of $128$. The length $Q$ of MTQs for a single task is set to $8$. We optimize both phases with the AdamW optimizer Kingma & Ba (2014). We implement all experiments on 8 NVIDIA H20 GPUs, each with $98$ GB of memory. We carry out the majority of metric evaluations using the VLEMEval library Duan et al. (2024), except when certain metrics are not covered. For detailed evaluation procedures, please refer to the Appendix.
@@ 4.1 Quantitative Comparison
Results on General Benchmarks. To assess the general capabilities of our VaCo, we conduct a comprehensive comparison with prominent MLLMs, as presented in Table 3.3. The benchmarks encompass diverse sub-tasks that evaluate various fine-grained capabilities, such as logic/attribute/relation understanding and coarse/fine-grained perception in MMBench Liu et al. (2024b). Benefiting from the stronger visual comprehension ability provided by vision-centric activation and coordination, VaCo achieves superior performance on these benchmarks. For instance, our VaCo attains an overall score of $78.6$ on MMBench-English and $46.7$ on MMMU Yue et al. (2024), outperforming LLaVA-1.5 Liu et al. (2023a) by $2.3$ points and $4.9$ points, respectively. Notably, compared to ROSS Wang et al. (2024a) that employs intrinsic visual activation with a reconstructive objective, our VaCo demonstrates significant advantages on benchmarks such as SEED Li et al. (2023a), MMVP Tong et al. (2024b), RealWorld, and AI2D Hiippala et al. (2021). This indicates that incorporating a query-based discriminative distillation objective with perceptual priors across multiple visual foundation models is more effective. Besides, our VaCo allows MLLMs to utilize a single visual encoder during inference, avoiding additional computational overhead from multiple encoders. Despite this, VaCo achieves superior performance across various metrics compared to Cambrian-1 Tong et al. (2024a), which aggregates multiple visual encoders (CLIP Radford et al. (2021), SigLIP Zhai et al. (2023), DINOv2 Oquab et al. (2023), and ConvNext Liu et al. (2022)). These results highlight the efficacy of vision-centric activation and coordination in improving the general visual-language understanding capabilities of MLLMs.
@@ 4.2 Ablation Study
In Table 4, we investigate the effects of various visual activation strategies and VFMs. We adopt LLaVA-v1.5 Liu et al. (2023a) baseline for all ablation studies, which integrates Vicuna-7B as the backbone and CLIP-ViT-L/14@336 as the vision encoder. We train our VaCo on LLavA-558K and Cambrian-737K datasets during the pre-training and the instruction tuning stages, respectively. All variants of VaCo are evaluated on (1) HaluBench Guan et al. (2024): Hall.; (2) MMBEN: MMBench-English Liu et al. (2024b); (3) SEEDI: SEED-Image Li et al. (2023a); (4) MMMU Yue et al. (2024); (5) MMVP Tong et al. (2024b); and (6) Real.: RealWorldQA x.ai (2024) benchmarks.
@@ 5.1 Multimodal Large Language Models
Recent advances in LLMs Achiam et al. (2023); Chiang et al. (2023); Touvron et al. (2023); Zeng et al. (2022) have driven substantial progress in multimodal LLMs (MLLMs) for visual recognition and understanding. LLM-based multimodal models Liu et al. (2023a); Bai et al. (2023); Chen et al. (2023a; 2024); Wang et al. (2023b); Zhai et al. (2022); Zhu et al. (2023) aim to integrate the advanced understanding capabilities of LLMs with the visual information provided by language-supervised visual encoders Radford et al. (2021); Zellers et al. (2019); Zhai et al. (2023); Tong et al. (2024b). A connector Liu et al. (2023a; 2024a); Alayrac et al. (2022); Li et al. (2023b); Ge et al. (2023); Li et al. (2024a) is employed to convert visual signals from the visual encoder into visual tokens within the LLM representational space. Despite these models exhibiting powerful performance, the lack of explicit visual supervision inherently leads to the degradation of rich visual information. Recent studies Wang et al. (2024a); Huang et al. (2025) have emphasized that activating visual signals boosts the comprehension capabilities of MLLMs. In contrast to ROSS Wang et al. (2024a), which reconstructs single visual features from MLLM visual tokens, our VaCo adopts a query-based pipeline that enables more flexible alignment of task-aware perceptual priors across multiple VFMs.
@@ 5.2 Vision Foundation Models in MLLMs
The pioneering LLaVA Liu et al. (2023a) model adopts CLIP Radford et al. (2021) as its visual encoder, and subsequent work has integrated stronger vision–language alignment encoders Zhai et al. (2023); Chen et al. (2024) into MLLM architectures. Building upon these approaches, some models Tong et al. (2024a; b) incorporate multiple visual encoders, such as DINOv2 Oquab et al. (2023), aMAE He et al. (2022) and MoCoV3 He et al. (2020), alongside vision-language alignment models. In recent years, certain works Jain et al. (2024); Huang et al. (2025) have discovered that visual signals from task-specific models address the limitation of CLIP, which mainly captures global information. Consequently, task-specific models such as DepthAnything Yang et al. (2024a; b), OneFormer Jain et al. (2023), FLARE Zhang et al. (2025) and VGGT Wang et al. (2025) have potential for comprehensive integration into MLLMs. Nonetheless, multi-encoder schemes, such as Cambrain-1 Tong et al. (2024a) and VCoder Jain et al. (2024), directly feed multi-source features into MLLMs, increasing the visual encoding cost while leaving potential representation conflicts unaddressed. Unlike these works, our VaCo leverages learnable task-specific queries to activate and coordinate the knowledge from multiple VFMs, thereby enhancing the visual comprehension capabilities of MLLMs without significantly increasing computational overhead.